### v2

@@ 4.6 Visual Turing Tests
Figure 6: The above are sample questions for a user survey. Each question presents four options, one of which is an image generated by this study. The questions evaluate the images’ overall aesthetic appeal, style consistency, creativity, and detail quality.
In our study, we used the same canny and style reference images to generate output with different models. We selected the best options based on four evaluation metrics: aesthetic appeal, style consistency, creativity, and detail quality, as shown in Figure 6. The results in Table 3 from these evaluations provide insights into the effectiveness of the LPen model in generating high-quality, artistically compelling landscape paintings.
@@
Table 3: This table presents the statistical data from the user survey questions. Each data point represents the top-rated results, as determined by users, for images generated by different models in terms of aesthetic appeal, style consistency, creativity, and detail quality.
{adjustwidth}
[…]
@@
In terms of style consistency, LPGen excelled with an impressive score of 61. 46%, far exceeding the Reference Only model (11.46%), the Double ControlNet model (16.66%) and the Lora model (10.42%). This high score indicates that LPGen-generated images consistently adhered to the specific artistic style more effectively, ensuring uniformity in stylistic elements crucial for cohesive and authentic artwork. When it comes to evaluating creativity, LPGen once again led the pack with a top score of 40.63%. This metric highlights the model’s ability to produce novel and inventive compositions within the realm of traditional landscape painting. In contrast, the Reference Only model scored 22.92%, the Double ControlNet model 23.95%, and the Lora model 12.50%. LPGen’s higher score in creativity underscores its capacity to generate images that are not only original but also creatively engaging. Regarding detail quality, LPGen achieved the highest score of 52.08%, showcasing its superiority in rendering intricate elements with precision and clarity. This capability is particularly important for realistically depicting natural features such as trees, rocks, and water. The Reference Only model scored 15.63%, the Double ControlNet model 19.79%, and the Lora model 12.50%. LPGen’s outstanding performance in detail quality highlights its proficiency in producing images with high fidelity and meticulous attention to fine details. Considering these results from all four metrics, the LPGen model’s outstanding performance across these metrics—aesthetic appeal, style consistency, creativity, and detail quality—highlights its effectiveness in generating high-quality, artistically compelling landscape paintings. These results reflect the model’s ability to meet and exceed user expectations in various aspects of image generation.
@@ 5 Conclusions
In the conclusion of this paper, we provide a comprehensive examination of our use of quantitative, qualitative, and user-study methodologies to evaluate the performance of various models in generating artistic-style images. Our research combined these approaches to offer a holistic assessment. Quantitative analysis used metrics such as aesthetic appeal, style consistency, creativity, and detail quality to measure model performance. For example, LPGen consistently outperformed other models in aesthetic appeal and detail quality, achieving the highest scores in these categories. Qualitative analysis delved deeper into user evaluations and perceptions of the generated images. This analysis revealed the strengths and weaknesses of different models in replicating artistic styles and expressing creativity. Through qualitative feedback, we gained insights into user experiences and preferences that quantitative metrics alone could not capture. User studies were pivotal in validating our quantitative and qualitative findings. These studies confirmed LPGen’s superiority in visual appeal and innovation from the user perspective, reinforcing the quantitative data and adding context to the qualitative insights.
In summary, our study employed a diverse set of research methodologies to thoroughly evaluate LPGen’s performance in generating artistic-style images. These analyses provided robust quantitative data and deep qualitative insights, enhancing our understanding and advancing the field of artistic generation models. We also noted that LPGen demands substantial computational resources, especially when processing large-scale data or generating high-resolution images. Future research will focus on enhancing the structure controller to more accurately manage generated structures, thereby further improving the quality and diversity of the generated images.
@@
The dataset used in this study will be made publicly available at a later stage and can be requested by contacting aiexplorercn@163.com.

### v3

@@ 4.4 Visual Assessment
Figure 6: Sample questions for a user survey. Each question presents four options, one of which is an image generated by this study. The questions evaSluate the images’ overall aesthetic appeal, style consistency, creativity, and detail quality.
A total of 24 groups of landscape paintings were generated using different models: LPGen, Reference Only, Double ControlNet, and Lora, for a questionnaire. Additionally, 16 artists were invited to participate in the survey. In this research, we evaluated four critical aspects of the generated images: aesthetic appeal, style consistency, creativity, and detail quality. The aesthetic appeal metric evaluates the overall visual attractiveness of the images. The participants rated the images based on how pleasing they found them, considering factors such as color harmony, composition, and the emotional response evoked by the artwork. The consistency aspect of the style examines how well the generated images adhere to a specific artistic style. Participants evaluated whether the images consistently incorporated stylistic elements of traditional landscape paintings, such as brushstroke techniques, use of space, and traditional motifs. The creativity aspect measures the originality and innovation of the generated images. The participants rated the images based on the novelty and inventiveness of the compositions and interpretations within the confines of traditional landscape painting. The detail quality metric focuses on the precision and clarity of the finer details within the images. Participants evaluated the quality of intricate elements such as textures, line work, and depiction of natural features such as trees, rocks, and water.
@@
Figure 7: Quantitative evaluation of the different models when generating landscape paintings. Each data represents the top-rated results, as determined by users, for images generated by different models in terms of aesthetic appeal, style consistency, creativity, and detail quality.
@@
From Figure 7, it can be seen that LPGen excelled with an impressive score of 61. 46%, far exceeding the Reference Only model, the Double ControlNet model, and the Lora model in terms of style consistency. When it comes to evaluating creativity, LPGen once again led the pack with a top score of 40.63%. In contrast, the Reference Only model scored 22.92%, the Double ControlNet model 23.95%, and the Lora model 12.50%. Regarding detail quality, LPGen achieved the highest score of 52.08%, showcasing its superiority in rendering intricate elements with precision and clarity. Moreover, LPGen has excellent aesthetic appearance, reaching the highest level of 57.28%. Considering these results from all four metrics, the LPGen model’s outstanding performance across four metrics—aesthetic appeal, style consistency, creativity, and detail quality—highlights its effectiveness in generating high-quality, artistically compelling landscape paintings. These results reflect the model’s ability to meet and exceed user expectations in various aspects of image generation.
4.5 Generated Showcase
Figure 8: Landscape paintings generated by the proposed method using the same structure reference image but different style images. The generated images have distinct stylistic features such as azure green landscape, ink wash landscape, and light vermilion landscape
Figure 8 shows the landscape paintings generated by the proposed method, which is in the style of the famous such as azure green landscape, ink wash landscape, and light vermilion landscape. The landscape paintings generated met the style requirements and were full of details. Specifically, the techniques in landscape painting such as outlining, chapping, rubbing, moss-dotting, and coloring are clearly visible, with distinct stylistic features that accurately capture the essence of traditional Chinese landscape painting. In addition, the three images initially used the same structure reference image. The generated images had a good composition, accurately distinguishing mountains, trees, and distant views. Therefore, the architectural design generated by the proposed method reached a usable level, capable of reducing creative difficulty and improving creative efficiency.
5 Discussion
This study demonstrates the effectiveness of the proposed method through both qualitative and quantitative analyses. In terms of qualitative analysis, visual comparisons with other methods show that the proposed method can generate landscape paintings with specified styles and structures, a capability that mainstream models lack. On the quantitative side, the data prove the superiority of the proposed method across all evaluation metrics. In particular, the method excels at generating specified styles and structures. For example, in the histogram similarity metric, the proposed method performs well, with values of 0.89, 0.81, and 0.75, respectively, lower than those of reference only, double controlNet, and Lora, thus convincingly demonstrating its effectiveness.
By directly utilizing AI to generate diverse landscape paintings, LPGen replaces the tedious processes of outlining, chapping, rubbing, moss-dotting, and coloring in the traditional hand-drawing method. Compared to conventional methods, LPGen excels in both efficiency and creative generation. Regarding design efficiency, traditional methods typically require approximately two days to complete a creation and its corresponding revisions, whereas LPGen, running on a consumer-grade GPU with 8GB VRAM, can generate a complex image description using Stable Diffusion in around 4 seconds. As computational power continues to improve, the speed of generating interior design videos with LPGen can be further enhanced. In terms of creative design, LPGen offers style options for users to choose from, allowing for either hand-drawn structure diagrams or automatic extraction of structure diagrams from other images, thereby accelerating the design process. In general, LPGen demonstrates the feasibility of innovative methods for generating landscape images. Additionally, LPGen is highly scalable; by replacing the underlying diffusion model, it can be adapted to other generation painting.
The content generated by AI will profoundly impact current painting methods. In terms of efficiency, AI will increasingly take over tasks that emphasize logical and rational descriptions, ultimately forming an AI design chain. Simple design tasks will be completed by AI, thereby giving designers more time to focus on creativity and improve quality. In terms of role positioning, designers are no longer just traditional creators; they are transitioning to facilitators collaborating with AI. For example, in this study, the artist’s role is not merely to draw images; they use their expertise to collect and organize data and transfer knowledge to the AI model. This new human- machine collaboration approach is likely to become the norm in future digital painting, driving the design process toward automation and intelligence.
@@ 6 Conclusions
Traditional hand-drawing methods for creating landscape paintings not only require the artist to have extensive experience and creativity but also involve labor-intensive efforts. This often leads to a lack of creativity and low design efficiency. To address these issues, we propose LPGen for the automatic generation of landscape paintings. Experimental results demonstrate that LPGen is a high-fidelity, controllable model capable of replacing the tedious creative and drawing tasks in the traditional design process. This transformation of the conventional landscape painting creation workflow significantly enhances design efficiency and creativity.
Nevertheless, this study acknowledges certain limitations that warrant further investigation. Firstly, controlling the generation of landscape paintings using canny edge maps and reference style images is challenging. Future improvements could be made by incorporating additional structural control methods such as depth maps, sketches, and recoloring. Secondly, while generating styles, the current method controls the style of the entire generated image, but it cannot specify the texture appearance in specific regions of the image. Therefore, enhancing the controllability of the generation process is necessary. Additionally, our understanding of painting styles is manually classified; a more extensive and automated annotation classification method might better represent real-world style classifications. Finally, the style metrics of the images generated by the proposed method can still be improved to enhance the realism and usability of the generated images.
AI can be applied to various stages of future painting creation. For example, in the conceptualization stage, AI can provide artists with a range of creative inspirations and style references to choose from. Subsequently, image generation technologies based on GANs or DMs can quickly produce numerous sketches and initial compositions. In the detail refinement stage, automated AI algorithms can optimize lines, colors, and textures, thereby reducing the workload of traditional manual retouching. Finally, automated style transfer and high-resolution rendering technologies can enhance the efficiency from creation to presentation, ultimately forming a more comprehensive AI-assisted painting creation workflow.
@@
The dataset used in this study will be made publicly available at a later stage and can be requested by contacting ywg@bigc.edu.cn.